OpenStudy (anonymous):

Write an equation in point-slope form, y – y1 = m(x – x1), of the line through points (9, –2) and (10, 2). Use (9, –2) as the point (x1, y1). a. (y + 2) = –4(x – 9) c. (y + 2) = 4(x – 9) b. (y – 2) = 4(x + 9) d. (y – 2) = –4(x + 9)

OpenStudy (anonymous):

Write an equation in point-slope form, y – y1 = m(x – x1), of the line through points (4, –4) and (1, 2). Use (4, –4) as the point (x1, y1).

OpenStudy (anonymous):

(y+4)=-2(x-4)